0.4(3(4)-2)+2(4)+4÷3
Answer by fractalier(6550)

Using PEMDAS, we get
0.4(3(4)-2) + 2(4) + 4÷3 =
0.4(12 - 2) + 8 + 4/3 =
0.4(10) + 8 + 4/3 =
4 + 8 + 4/3 =
13 1/3 or 40/3